# Sorbonne Université & CNRS, 4 place Jussieu, 75252 Paris Cedex 05, France

### v1

@@ 1 Introduction
The discovery of CP violation in kaon systems played a crucial role in motivating the existence of the top quark, highlighting the requirement for a third family of quarks to enable a CP-violating phase in the Cabibbo-Kobayashi-Maskawa (CKM) matrix. However, the experimental confirmation of the top quark was achieved only in 1995 by the CDF and D0 experiments at Fermilab CDF:1995wbb ; D0:1995jca . Thirty years later, the top quark is recognised as the heaviest among all known elementary particles with a mass that originates, in the Standard Model (SM), from the top interaction with the Higgs field and the related sizeable Yukawa coupling. This suggests that the top quark has a special role in electroweak symmetry breaking, and that it is possibly connected to yet-unknown new dynamics. It is therefore widely regarded as a powerful probe for physics beyond the SM (BSM). Consequently, measurements involving top quark production and decay processes, as well as precisely measured observables where the top contributes via quantum loops, are deeply scrutinised as they can potentially provide vital information about BSM scenarios Atwood:2000tu .
@@
With the ongoing third run of the LHC, both the CMS and ATLAS experiments will soon have access to about 500 fb-1 of data. This large integrated luminosity opens the possibility to search for rare phenomena corresponding to production cross sections of only a few fb. Among all these rare processes of potential interest, those depending on flavour-changing neutral current (FCNC) interactions involving the top quark are particularly relevant. In the SM, top FCNCs are highly suppressed due to the Glashow-Iliopoulos-Maiani (GIM) mechanism, leading to extremely small branching fractions for decays such as $t\to qZ$, $t\to qH$, $t\to q\gamma$ and $t\to qg$ with $q=u,c$, typically in the range $10^{-17}-10^{-12}$ Eilam:1990zc ; Mele:1998ag ; Aguilar-Saavedra:2002lwv ; Aguilar-Saavedra:2004mfd ; Zhang:2013xya ; Durieux:2014xla . While these rates lie far beyond the reach of current experiments, various extensions of the SM can enhance them by several orders of magnitude, sometimes up to $\mathcal{O}(10^{-7})-\mathcal{O}(10^{-4})$ Castro:2022qkg ; Durieux:2014xla ; Larios:2006pb ; Barros:2019wxe , bringing them within the expected sensitivity of the LHC experiments. Any observation of a deviation compatible with FCNC effects in the top sector would therefore constitute a strong indication of BSM physics associated with flavour, as typically expected, for instance, in two-Higgs-doublet models Atwood:1996vj ; Botella:2015hoa ; Abbas:2015cua ; Baum:2008qm , the minimal supersymmetric standard model Dedes:2014asa ; Cao:2007dk ; Lopez:1997xv ; Eilam:2001dh , left-right (super)symmetric Gaitan:2004by ; Frank:2005vd ; Frank:2023fkc models, extra dimensional models Gao:2013fxa ; Dey:2016cve ; Diaz-Furlong:2016ril and dark matter models Agrawal:2014aoa ; Agrawal:2015kje ; Jueid:2024cge .
@@
Top quark FCNCs are typically studied at colliders via two main analysis strategies: top pair production followed by a rare FCNC decay of one of the two produced top quarks, and FCNC single top production followed by a standard top decay. To date, no significant deviation from the SM expectations has been observed in searches performed at LEP L3:2002hbp ; OPAL:2001spi ; ALEPH:2002wad , HERA ZEUS:2003vfj ; ZEUS:2011mya and the Tevatron Kikuchi:2000sv ; D0:2010dry ; D0:2007wfn . At the LHC, both the ATLAS and CMS collaborations have carried out dedicated searches in multiple channels, and current bounds on branching ratios such as $\text{BR}(t\to qZ)$, $\text{BR}(t\to q\gamma)$, $\text{BR}(t\to qH)$ and $\text{BR}(t\to qg)$ are now typically in the range $10^{-5}-10^{-3}$ ATLAS:2023ujo ; ATLAS:2023qzr ; ATLAS:2022gzn ; ATLAS:2021amo ; ATLAS:2018zsq ; ATLAS:2018jqi ; ATLAS:2018xxe ; CMS:2016uzc ; CMS:2017wcz ; CMS:2021hug ; CMS:2020utv ; ATLAS:2024mih ; CMS:2024ubt ; ATLAS:2015vhj ; CMS:2013knb ; CMS:2015kek ; ATLAS:2022per ; CMS:2023bjm ; ATLAS:2015iqc ; ATLAS:2012lhi . While they lie still several orders of magnitude above SM expectations, these results can already be used to constrain a variety of new physics models. Furthermore, with the high-luminosity phase of the LHC (HL-LHC) at the horizon, sensitivity to these rare decay processes is expected to significantly improve due to increased luminosity and better detector performance, thus potentially probing (or even ruling out) broad classes of BSM scenarios. In light of the anticipated HL-LHC results on top FCNCs, it hence becomes timely to revisit the analysis strategies used to search for these rare processes. In this work, we investigate the potential of modern collider analysis methodologies to boost sensitivity to FCNC signals in top quark events. Our goal is to assess the added value of deep learning approaches over traditional methods by systematically comparing a conventional cut-based analysis to several deep learning strategies that we apply to a set of low-level and high-level reconstructed kinematic variables. Deep learning methods often outperform traditional cut-based approaches due to their ability to learn complex and non-linear decision boundaries in high-dimensional feature spaces. While cut-based analyses rely on manually-defined thresholds applied to a limited set of kinematic variables, thus potentially overlooking important correlations, deep learning techniques can automatically extract and integrate information from all input features. This subsequently enables them to identify subtle patterns that can be used to distinguish signal from background, resulting in significantly improved classification performance. In particular, we consider three distinct deep learning architectures that have shown promising performance in recent studies: multi-layer perceptrons (MLPs) Rumelhart:1986 ; Goodfellow:2016 , graph attention networks (GATs) Velickovic:2017lzs ; Yu:2023juh ; Calafiura:2024qhv ; Esmail:2023axd and transformer-based models Qu:2022mxj ; Wu:2024thh ; Hammad:2023sbd ; Hammad:2024cae . These architectures differ in how they handle the input data structure and exploit correlations among observables, offering a broad perspective on the role of representation learning in collider physics. GATs and transformer architectures offer significant advantages over MLPs in classification tasks, especially when handling structured or relational data. GATs are particularly well-suited for capturing interactions between reconstructed final-state particles within graph-based event representations, effectively modelling complex topologies and local connectivity patterns that MLPs are not inherently designed to process. Transformers, by contrast, employ self-attention mechanisms to dynamically evaluate the importance of each input particle, assigning context-dependent weights that help distinguish between signal and background. Unlike MLPs, which treat input features as fixed-size unstructured vectors, both GATs and transformers can adapt to the underlying event structure, yielding more expressive models and enhanced classification performance. By comparing their performance across realistic simulated FCNC signal and background samples, we aim to highlight both the strengths and limitations of these approaches and identify possible paths forward for future experimental searches at the LHC.
@@
To provide quantitative estimates of the potential improvements in sensitivity, we work within the framework of the SM effective field theory (SMEFT) which offers a model-independent parametrisation of new physics effects via higher-dimensional operators that respect the SM symmetries Buchmuller:1985jz ; Grzadkowski:2010es ; Brivio:2017vri . This bottom-up approach is particularly well motivated when possible new particles lie at scales beyond the current experimental reach, and then allows for a consistent and unified treatment of top-quark FCNC processes in both production and decay channels. For practical purposes and in order to streamline the analysis, we describe the relevant BSM interactions in terms of a minimal set of effective operators defined in the broken electroweak phase Aguilar-Saavedra:2008nuh . Furthermore, as a case study we concentrate on processes involving an anomalous coupling between the top quark, a photon and a light up-type quark ($u$ or $c$), and we simulate both the related single-top and top-antitop production channels at the LHC. The performance of the different explored analysis strategies is next evaluated and compared in terms of sensitivity to the FCNC signal considered for a luminosity of 500 fb-1. Moreover, we also confront our predictions with existing LHC Run 2 data, normalising our results to an integrated luminosity of 139 fb-1 to better assess quantitatively the potential impact of the proposed methods. Finally, although our analysis is applicable to other top FCNC couplings, different collider luminosities and centre-of-mass energies, the associated comprehensive exploration is deferred to future work.
@@
The rest of this report is organised as follows. In section 2, we present our effective parametrisation of top-quark FCNC interactions and summarise the current experimental constraints from a broad range of experiments. In section 3, we introduce our simulation framework and provide details on the generation of both background and signal events in the case of a top FCNC interaction with the photon. Section 4 outlines the various analysis strategies employed in this study, including both a conventional cut-based approach and three different machine learning methods. In section 5, we report our results for top FCNC interactions involving a photon and a light up-type quark at LHC luminosities of 139 and 500 fb-1. We finally summarise our findings and conclude in section 6.
@@ 2 Top-quark FCNC interactions at colliders: Lagrangians and bounds
At high-energy colliders such as the LHC, searches for top-quark FCNC interactions can be pursued through two complementary approaches, namely the study of rare top decays in top-antitop pair production events (for example from $pp\to t\bar{t}\to qX\,Wb$ where $X$ denotes a SM gauge or Higgs boson) and the investigation of FCNC-induced single top production processes (like $pp\to tX$). In this work, we consider the joint effect of these two channels as probes to non-standard top-quark couplings to lighter quarks and neutral bosons. In addition, to interpret potential deviations from the SM predictions in a consistent and model-independent way, we adopt the SMEFT framework Buchmuller:1985jz ; Grzadkowski:2010es ; Brivio:2017vri where the SM Lagrangian ${\cal{L}}_{\mathrm{SM}}$ is extended by higher-dimensional operators constructed from the SM fields and respecting the full $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ gauge symmetry,
@@
In the case of top-quark FCNC interactions, several dimension-six SMEFT operators contribute to both top decay and production processes. These include operators modifying the top’s interactions with the SM gauge bosons and light quarks, as well as four-fermion interactions involving at least one top-quark field. The relevant operators for this work are of the first category, and they can be grouped, following the normalisation convention and naming scheme of Zhang:2014rja , as
$\begin{split}\mathcal{O}_{\varphi q}^{1(ij)}=\frac{y_{t}^{2}}{2}\,(\varphi^{\dagger}i\overleftrightarrow{D}_{\!\!\mu}\varphi)\,(\bar{q}_{i}\gamma^{\mu}q_{j})\,,\qquad\qquad&\mathcal{O}_{\varphi q}^{3(ij)}=\frac{y_{t}^{2}}{2}\,(\varphi^{\dagger}i\overleftrightarrow{D}_{\!\!\mu}^{\!I}\varphi)\,(\bar{q}_{i}\gamma^{\mu}\tau^{I}q_{j})\,,\\ \mathcal{O}_{\varphi u}^{(ij)}=\frac{y_{t}^{2}}{2}\,(\varphi^{\dagger}i\overleftrightarrow{D_{\mu}}\varphi)\,(\bar{u}_{i}\gamma^{\mu}u_{j})\,,\qquad\qquad&\mathcal{O}_{u\varphi}^{(ij)}=-y_{t}^{3}\,(\bar{q}_{i}u_{j}\tilde{\varphi})\,(\varphi^{\dagger}\varphi)\,,\\ \mathcal{O}_{uW}^{(ij)}=y_{t}g\,(\bar{q}_{i}\sigma^{\mu\nu}\tau^{I}u_{j})\,\tilde{\varphi}W_{\mu\nu}^{I}\,,\qquad\qquad&\mathcal{O}_{uB}^{(ij)}=y_{t}g_{Y}\,(\bar{q}_{i}\sigma^{\mu\nu}u_{j})\,\tilde{\varphi}B_{\mu\nu}\,,\\[4.26773pt] \mathcal{O}_{uG}^{(ij)}=y_{t}g_{s}\,(\bar{q}_{i}\sigma^{\mu\nu}T^{A}u_{j})\,\tilde{\varphi}G_{\mu\nu}^{A}\,.\qquad\qquad&\end{split}$ (2)
In these expressions, $q_{i}$ and $u_{i}$ denote the left-handed quark doublet and right-handed up-type quark singlet fields of generation $i$, respectively, while $\varphi$ represents the SM Higgs doublet (with $\tilde{\varphi}=i\tau^{2}\varphi^{*}$). Moreover, $B_{\mu\nu}$, $W_{\mu\nu}^{I}$ and $G_{\mu\nu}^{A}$ are the field strength tensors of the $U(1)_{Y}$, $SU(2)_{L}$ and $SU(3)_{C}$ gauge fields, with the corresponding gauge couplings being given by $g_{Y}$, $g$ and $g_{s}$. In addition, we denote by $\tau^{I}$ and $T^{A}$ the $SU(2)$ and $SU(3)$ generators in the fundamental representation, respectively, and by $Y$ the hypercharge operator. Finally, the covariant derivative operator acting on the Higgs field is defined as usual, with $D_{\mu}=\partial_{\mu}-ig\tau^{I}W_{\mu}^{I}-ig_{Y}YB_{\mu}$, and the bidirectional derivatives are given by
@@
We also emphasise that according to the normalisation convention of Zhang:2014rja , explicit powers of the top Yukawa coupling $y_{t}$ are factored out in the operator definitions. Among the electroweak operators $\mathcal{O}_{\varphi q}^{1}$ and $\mathcal{O}_{\varphi q}^{3}$, only the linear combination $\mathcal{O}_{\varphi q}^{-(ij)}\equiv\mathcal{O}_{\varphi q}^{1(ij)}-\mathcal{O}_{\varphi q}^{3(ij)}$ contributes to up-sector FCNC transitions at tree level, with the orthogonal combination $\mathcal{O}_{\varphi q}^{+}$ defined in a similar way affecting instead down-type quark interactions. Restricting our analysis to flavour-changing interactions between the top quark and the first two generations of light up-type quarks, we therefore define a BSM parameter space built on a set of ten independent complex parameters per light quark flavour $a=1,2$ Durieux:2014xla ,
@@
In this Lagrangian, $q=u,c$ denotes again a light up-type quark, while the effective couplings $v_{tq}^{Z}$, $a_{tq}^{Z}$, $g_{qt}^{v,a}$, $f_{tq}^{V}$ and $h_{tq}^{V}$ (with $V=\gamma,Z,g$) parametrise the strength of the FCNC interactions of the top quark with the Higgs boson $h$, the $Z$ boson, the photon $A$ and the gluon $g$, respectively, with the corresponding gauge boson field-strength tensors being represented by $A_{\mu\nu}$, $Z_{\mu\nu}$ and $G_{\mu\nu}^{A}$. Moreover, we conventionally introduced in the prefactors of the top FCNC couplings to the $Z$ boson a dependence on the cosine of the electroweak mixing angle $c_{W}$. This broken-phase parametrisation features only two dimension-four operators (the vector and scalar currents) and three dimension-five dipole operators per flavour. While it does not reflect the full gauge structure of SMEFT at high energies, it provides, as said above, a convenient and widely used framework for modelling FCNC processes in the top-quark sector Aguilar-Saavedra:2004mfd ; Aguilar-Saavedra:2008nuh ; Durieux:2014xla .
@@
Our goal is then to assess the sensitivity of analysis strategies relying on either a traditional cut-and-count approach or deep learning methods, and to interpret the results in terms of bounds on the effective dipole couplings $f_{tq}^{\gamma}$ and $h_{tq}^{\gamma}$ introduced in Eq. (5). However, experimental constraints on top FCNC interactions are commonly reported in terms of upper bounds on the top rare branching ratios. We therefore provide expressions for the partial width of the decay $t\to q\gamma$ in terms of the dipole coefficients Beneke:2000hk ; Aguilar-Saavedra:2004mfd ,
@@
where $\alpha=e^{2}/(4\pi)$ is the electromagnetic coupling constant and $m_{t}$ is the top-quark mass. To extract the associated branching ratio, we normalise this result to the total width of the top quark that includes, in addition to Eq. (7), the dominant contribution from the SM process $t\to bW^{+}$. At next-to-leading order in the strong coupling $\alpha_{s}=g_{s}^{2}/(4\pi)$, the latter is given by Beneke:2000hk
@@
Over the past decade, both the CMS and ATLAS collaborations have conducted dedicated searches for FCNC interactions involving the top quark and a photon. The first constraints on the $\mathrm{BR}(t\to q\gamma)$ branching ratio of Eq. (9) came from CMS, using the full 8 TeV dataset to set limits of BR$(t\rightarrow u\gamma)<1.3\times 10^{-4}$ and BR$(t\rightarrow c\gamma)<1.7\times 10^{-3}$ at 95% confidence level (CL) CMS:2015kek . More recently, ATLAS improved these bounds using 139 fb-1 of 13 TeV data, reporting BR$(t\rightarrow u\gamma)<(0.85-1.2)\times 10^{-5}$ and BR$(t\rightarrow c\gamma)<(4.2-4.5)\times 10^{-5}$ at 95% CL ATLAS:2022per , the precise threshold depending on the Lorentz structure of the top FCNC interactions. However, the latest and most stringent constraints have been established by CMS using the full Run 2 dataset CMS:2023bjm . This last analysis combined events originating from single top production with an associated photon with $t\bar{t}$ events featuring a rare top decay. It then employed multivariate techniques to efficiently suppress the SM background, obtaining 95% CL upper limits of BR$(t\rightarrow\gamma u)<0.95\times 10^{-5}$ and BR$(t\rightarrow\gamma c)<1.51\times 10^{-5}$.
Looking ahead, the HL-LHC is expected to significantly improve the sensitivity to rare top decays. Operating at $\sqrt{s}=14$ TeV and aiming for an integrated luminosity of 3000 fb-1, projections suggest that branching fractions as low as $10^{-6}$ could be probed Mandrik:2018gud . Additionally, even more stringent limits are anticipated from next-generation colliders. A future circular hadron collider foreseen to operate at a centre-of-mass energy of $\sqrt{s}=100$ TeV with 30 ab-1 of data could hence push the sensitivity down to BR$(t\rightarrow u\gamma)<1.8\times 10^{-7}$ and BR$(t\rightarrow c\gamma)<2.4\times 10^{-7}$ Mandrik:2018yhe , while other proposed facilities such as TESLA Aguilar-Saavedra:2001ajk and future muon colliders Ake:2023xcz could in principle provide complementary information due to their clean experimental environments and potential for polarisation control.

### v2

@@ 1 Introduction
The discovery of CP violation in kaon systems played a crucial role in motivating the existence of the top quark, highlighting the requirement for a third family of quarks to enable a CP-violating phase in the Cabibbo-Kobayashi-Maskawa (CKM) matrix. However, the experimental confirmation of the top quark was achieved only in 1995 by the CDF and D0 experiments at Fermilab Abe and others (1995); Abachi and others (1995). Thirty years later, the top quark is recognised as the heaviest among all known elementary particles with a mass that originates, in the Standard Model (SM), from the top interaction with the Higgs field and the related sizeable Yukawa coupling. This suggests that the top quark has a special role in electroweak symmetry breaking, and that it is possibly connected to yet-unknown new dynamics. It is therefore widely regarded as a powerful probe for physics beyond the SM (BSM). Consequently, measurements involving top quark production and decay processes, as well as precisely measured observables where the top contributes via quantum loops, are deeply scrutinised as they can potentially provide vital information about BSM scenarios Atwood et al. (2001).
@@
With the ongoing third run of the LHC, both the CMS and ATLAS experiments will soon have access to about 500 fb-1 of data. This large integrated luminosity opens the possibility to search for rare phenomena corresponding to production cross sections of only a few fb. Among all these rare processes of potential interest, those depending on flavour-changing neutral current (FCNC) interactions involving the top quark are particularly relevant. In the SM, top FCNCs are highly suppressed due to the Glashow-Iliopoulos-Maiani (GIM) mechanism, leading to extremely small branching fractions for decays such as $t\to qZ$, $t\to qH$, $t\to q\gamma$ and $t\to qg$ with $q=u,c$, typically in the range $10^{-17}-10^{-12}$ Eilam et al. (1991); Mele et al. (1998); Aguilar-Saavedra and Nobre (2003); Aguilar-Saavedra (2004); Zhang and Maltoni (2013); Durieux et al. (2015). While these rates lie far beyond the reach of current experiments, various extensions of the SM can enhance them by several orders of magnitude, sometimes up to $\mathcal{O}(10^{-7})-\mathcal{O}(10^{-4})$ Castro and Skovpen (2022); Durieux et al. (2015); Larios et al. (2006); Barros et al. (2020), bringing them within the expected sensitivity of the LHC experiments. Any observation of a deviation compatible with FCNC effects in the top sector would therefore constitute a strong indication of BSM physics associated with flavour, as typically expected, for instance, in two-Higgs-doublet models Atwood et al. (1997); Botella et al. (2016); Abbas et al. (2015); Baum et al. (2008); Balaji (2020), the minimal supersymmetric standard model Dedes et al. (2014); Cao et al. (2007); Lopez et al. (1997); Eilam et al. (2001), left-right (super)symmetric Gaitan et al. (2005); Frank and Turan (2005); Frank et al. (2024a) models, extra dimensional models Gao et al. (2013); Dey and Jha (2016); Diaz-Furlong et al. (2016), dark matter Agrawal et al. (2014, 2016); Jueid and Kanemura (2024) and composite models Balaji (2022); Crivellin et al. (2022).
Top quark FCNCs are typically studied at colliders via two main analysis strategies: top pair production followed by a rare FCNC decay of one of the two produced top quarks, and FCNC single top production followed by a standard top decay. To date, no significant deviation from the SM expectations has been observed in searches performed at LEP Achard and others (2002); Abbiendi and others (2001); Heister and others (2002), HERA Chekanov and others (2003); Abramowicz and others (2012) and the Tevatron Kikuchi et al. (2001); Abazov and others (2010, 2007). At the LHC, both the ATLAS and CMS collaborations have carried out dedicated searches in multiple channels, and current bounds on branching ratios such as $\text{BR}(t\to qZ)$, $\text{BR}(t\to q\gamma)$, $\text{BR}(t\to qH)$ and $\text{BR}(t\to qg)$ are now typically in the range $10^{-5}-10^{-3}$ Aad and others (2023c, a, b, 2022); Aaboud and others (2018a, 2019, b); Khachatryan and others (2017); Sirunyan and others (2017); Tumasyan and others (2022a); Sirunyan and others (2021); Aad and others (2024); Hayrapetyan and others (2024b); Aad and others (2016a); Chatrchyan and others (2014); Khachatryan and others (2016); Aad and others (2023d); Hayrapetyan and others (2024a); Aad and others (2016b, 2012). While they lie still several orders of magnitude above SM expectations, these results can already be used to constrain a variety of new physics models. Furthermore, with the high-luminosity phase of the LHC (HL-LHC) at the horizon, sensitivity to these rare decay processes is expected to significantly improve due to increased luminosity and better detector performance, thus potentially probing (or even ruling out) broad classes of BSM scenarios. In light of the anticipated HL-LHC results on top FCNCs, it hence becomes timely to revisit the analysis strategies used to search for these rare processes.
@@
In this work, we investigate the potential of modern collider analysis methodologies to boost sensitivity to FCNC signals in top quark events. Our goal is to compare several deep learning strategies that we apply to a set of low-level and high-level reconstructed kinematic variables. Deep learning methods often outperform traditional cut-based approaches due to their ability to learn complex and non-linear decision boundaries in high-dimensional feature spaces. This subsequently enables them to identify subtle patterns that can be used to distinguish signal from background, resulting in significantly improved classification performance. We consider three distinct deep learning architectures that have shown promising performance in recent studies: multi-layer perceptrons (MLPs) Rumelhart et al. (1986); Goodfellow et al. (2016), graph attention networks (GATs) Veličković et al. (2017); Yu et al. (2023); Calafiura et al. (2024); Esmail et al. (2023) and transformer-based models Qu et al. (2022); Wu et al. (2025); Hammad et al. (2024b); Hammad and Nojiri (2024). These architectures differ in how they handle the input data structure and exploit correlations among observables, offering a broad perspective on the role of representation learning in collider physics. GATs and transformer architectures offer significant advantages over MLPs in classification tasks, especially when handling structured or relational data. GATs are particularly well-suited for capturing interactions between reconstructed final-state particles within graph-based event representations, effectively modelling complex topologies and local connectivity patterns that MLPs are not inherently designed to process. Transformers, by contrast, employ self-attention mechanisms to dynamically evaluate the importance of each input particle, assigning context-dependent weights that help distinguish between signal and background. Unlike MLPs, which treat input features as fixed-size unstructured vectors, both GATs and transformers can adapt to the underlying event structure, yielding more expressive models and enhanced classification performance. By comparing their performance across realistic simulated FCNC signal and background samples, we aim to highlight both the strengths and limitations of these approaches and identify possible paths forward for future experimental searches at the LHC.
@@
To provide quantitative estimates of the potential improvements in sensitivity, we work within the framework of the SM effective field theory (SMEFT) which offers a model-independent parametrisation of new physics effects via higher-dimensional operators that respect the SM symmetries Buchmuller and Wyler (1986); Grzadkowski et al. (2010); Brivio and Trott (2019). This bottom-up approach is particularly well motivated when possible new particles lie at scales beyond the current experimental reach, and then allows for a consistent and unified treatment of top-quark FCNC processes in both production and decay channels. For practical purposes and in order to streamline the analysis, we describe the relevant BSM interactions in terms of a minimal set of effective operators defined in the broken electroweak phase Aguilar-Saavedra (2009). Furthermore, as a case study we concentrate on processes involving an anomalous coupling between the top quark, a photon and a light up-type quark ($u$ or $c$), and we simulate both the related single-top and top-antitop production channels at the LHC. The performance of the different explored analysis strategies is next evaluated and compared in terms of sensitivity to the FCNC signal considered for a luminosity of 500 fb-1. Moreover, we also confront our predictions with existing LHC Run 2 data, normalising our results to an integrated luminosity of 139 fb-1 to better assess quantitatively the potential impact of the proposed methods. Finally, although our analysis is applicable to other top FCNC couplings, different collider luminosities and centre-of-mass energies, the associated comprehensive exploration is deferred to future work.
@@
The rest of this report is organised as follows. In section 2, we present our effective parametrisation of top-quark FCNC interactions and summarise the current experimental constraints from a broad range of experiments. In section 3, we introduce our simulation framework and provide details on the generation of both background and signal events in the case of a top FCNC interaction with the photon. Section 4 outlines the various analysis strategies employed in this study. In section 5, we report our results for top FCNC interactions involving a photon and a light up-type quark at LHC luminosities of 139 and 500 fb-1. We finally summarise our findings and conclude in section 6.
@@ 2 Top-quark FCNC interactions at colliders: Lagrangians and bounds
At high-energy colliders such as the LHC, searches for top-quark FCNC interactions can be pursued through two complementary approaches, namely the study of rare top decays in top-antitop pair production events (for example from $pp\to t\bar{t}\to qX\,Wb$ where $X$ denotes a SM gauge or Higgs boson) and the investigation of FCNC-induced single top production processes (like $pp\to tX$). In this work, we consider the joint effect of these two channels as probes to non-standard top-quark couplings to lighter quarks and neutral bosons. In addition, to interpret potential deviations from the SM predictions in a consistent and model-independent way, we adopt the SMEFT framework Buchmuller and Wyler (1986); Grzadkowski et al. (2010); Brivio and Trott (2019) where the SM Lagrangian ${\cal{L}}_{\mathrm{SM}}$ is extended by higher-dimensional operators constructed from the SM fields and respecting the full $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ gauge symmetry,
@@
In the case of top-quark FCNC interactions, several dimension-six SMEFT operators contribute to both top decay and production processes. These include operators modifying the top’s interactions with the SM gauge bosons and light quarks, as well as four-fermion interactions involving at least one top-quark field. The relevant operators for this work are of the first category, and they can be grouped, following the normalisation convention and naming scheme of Zhang (2014), as
[…]
In these expressions, $q_{i}$ and $u_{i}$ denote the left-handed quark doublet and right-handed up-type quark singlet fields of generation $i$, respectively, while $\varphi$ represents the SM Higgs doublet (with $\tilde{\varphi}=i\tau^{2}\varphi^{*}$). Moreover, $B_{\mu\nu}$, $W_{\mu\nu}^{I}$ and $G_{\mu\nu}^{A}$ are the field strength tensors of the $U(1)_{Y}$, $SU(2)_{L}$ and $SU(3)_{C}$ gauge fields, with the corresponding gauge couplings being given by $_{Y}, gg$ and $g_{s}$. In addition, we denote by $\tau^{I}$ and $T^{A}$ the $SU(2)$ and $SU(3)$ generators in the fundamental representation, respectively, and by $Y$ the hypercharge operator. Finally, the covariant derivative operator acting on the Higgs field is defined as usual, with $D_{\mu}=\partial_{\mu}-ig\tau^{I}W_{\mu}^{I}-ig_{Y}YB_{\mu}$, and the bidirectional derivatives are given by
@@
We also emphasise that according to the normalisation convention of Zhang (2014), explicit powers of the top Yukawa coupling $y_{t}$ are factored out in the operator definitions. Among the electroweak operators $\mathcal{O}_{\varphi q}^{1}$ and $\mathcal{O}_{\varphi q}^{3}$, only the linear combination $\mathcal{O}_{\varphi q}^{-(ij)}\equiv\mathcal{O}_{\varphi q}^{1(ij)}-\mathcal{O}_{\varphi q}^{3(ij)}$ contributes to up-sector FCNC transitions at tree level, with the orthogonal combination $\mathcal{O}_{\varphi q}^{+}$ defined in a similar way affecting instead down-type quark interactions. Restricting our analysis to flavour-changing interactions between the top quark and the first two generations of light up-type quarks, we therefore define a BSM parameter space built on a set of ten independent complex parameters per light quark flavour $a=1,2$ Durieux et al. (2015),
@@
In this Lagrangian, $q=u,c$ denotes again a light up-type quark, while the effective couplings $v_{tq}^{Z}$, $a_{tq}^{Z}$, $g_{qt}^{v,a}$, $f_{tq}^{V}$ and $h_{tq}^{V}$ (with $V=\gamma,Z,g$) parametrise the strength of the FCNC interactions of the top quark with the Higgs boson $h$, the $Z$ boson, the photon $A$ and the gluon $g$, respectively, with the corresponding gauge boson field-strength tensors being represented by $A_{\mu\nu}$, $Z_{\mu\nu}$ and $G_{\mu\nu}^{A}$. Moreover, we conventionally introduced in the prefactors of the top FCNC couplings to the $Z$ boson a dependence on the cosine of the electroweak mixing angle $c_{W}$. This broken-phase parametrisation features only two dimension-four operators (the vector and scalar currents) and three dimension-five dipole operators per flavour. While it does not reflect the full gauge structure of SMEFT at high energies, it provides, as said above, a convenient and widely used framework for modelling FCNC processes in the top-quark sector Aguilar-Saavedra (2004, 2009); Durieux et al. (2015).
@@
We will then assess the sensitivity of analysis strategies relying on either a traditional cut-and-count approach or deep learning methods, and to interpret the results in terms of bounds on the effective dipole couplings $f_{tq}^{\gamma}$ and $h_{tq}^{\gamma}$ introduced in Eq. (5). However, experimental constraints on top FCNC interactions are commonly reported in terms of upper bounds on the top rare branching ratios. We therefore provide expressions for the partial width of the decay $t\to q\gamma$ in terms of the dipole coefficients Beneke and others (2000); Aguilar-Saavedra (2004),
@@
where $\alpha=e^{2}/(4\pi)$ is the electromagnetic coupling constant and $m_{t}$ is the top-quark mass. To extract the associated branching ratio, we normalise this result to the total width of the top quark that includes, in addition to Eq. (7), the dominant contribution from the SM process $t\to bW^{+}$. At next-to-leading order in the strong coupling $\alpha_{s}=g_{s}^{2}/(4\pi)$, the latter is given by Beneke and others (2000)
@@
[…]